OpenStudy (anonymous):

Please help me solve this word problem: In a round-robin chess tournament, each player is paired with every other player once. The function shown below, models the number of chess games, "N", that must be played in a round-robin tournament with "t" chess players. In a round-robin chess tournament, 55 games were played. How many players entered the tournament? N=t^2-t divided by 2.

OpenStudy (anonymous):

solve \[55=\frac{t^2-t}{2}\] for t

OpenStudy (anonymous):

\[t^2-t=110\] \[t^2-t-110=0\] \[(t-11)(t+10)=0\] \[t=11\]

OpenStudy (anonymous):

ahh thank you! I couldn't figure out haw to properly factor it